Suppose a particle P is moving in the plane so that its coordinates are given by P(x,y), where x = 4cos2t, y = 7sin2t.

(i) By finding a,b ∈ R such that x2 a2 + y2 b2 = 1, show that P is travelling on an elliptical path.

(ii) Let L(t) be the distance from P to the origin. Obtain an expression for L(t).

(iii) How fast is the distance between P and the origin changing when t = π/8?[


1
Expert's answer
2020-07-01T15:04:26-0400

"x = 4cos\ 2t; y = 7sin\ 2t\implies \frac{x}{4}=cos \ 2t...(1)""\frac{y}{7}=sin\ 2t...(2)"

(i) Squaring and adding equation (1) and (2), we get

"(\frac{x}{4})^2+(\frac{y}{7})^2=cos^2\ 2t+sin^2\ 2t=1"

So the equation becomes "(\frac{x^2}{4^2})+(\frac{y^2}{7^2})=1"

Hence,"a=4;b=7."

(ii) "L(t)=\sqrt{(x-0)^2+(y-0)^2}=\sqrt{x^2+y^2}"

"L(t)=\sqrt{(4cos\ 2t)^2+(7sin\ 2t)^2}" "=\sqrt{16cos^2\ 2t+49sin^2\ 2t}"

"L(t)=\sqrt{16+(49-16)sin^2 \ 2t}=\sqrt{16+33sin^2\ 2t}"

(iii) "\frac{dL}{dt}=" "\frac{d}{dt}(" "\sqrt{16+33sin^2\ 2t})" "=\frac{1}{2}\frac{1}{\sqrt{16+33sin^2\ 2t}}33\times 2sin\ 2t\ cos \ 2t\times 2" "=\frac{33}{\sqrt{16+33sin^2\ 2t}}sin\ 4t"

On putting "t=\frac{\pi}{8}" ,

"\frac{dL}{dt}=" "\frac{33}{\sqrt{16+33sin^2\ \frac{2\pi}{8}}}sin\ \frac{4\pi}{8}=" "\frac{33}{\sqrt{16+33sin^2\ \frac{\pi}{4}}}sin\ \frac{\pi}{2}=" "\frac{33}{\sqrt{16+33(\frac{1}{\sqrt{2}})^2}}=5.79"